# The non-Hermitian skin effect: A perspective

### v1

@@ 4.1 Beyond spectral winding numbers: Symmetries
There are models, which feature a NH skin effect, while having a zero spectral winding number. In these cases, it is possible to find alternative spectral invariants. For example, the presence of so-called time-reversal dagger symmetry (TRS[62] with the symmetry operator squaring to minus one results in doubly degenerate loops in the PBC spectrum with opposite winding numbers. As a result half of the skin states accumulate on one boundary and the other half on the other. The spectral invariant is a $\mathbb{Z}_{2}$ invariant, and this type of skin effect is referred to as the $\mathbb{Z}_{2}$-skin effect [28]. In general, symmetries play an important role in the context of the NH skin effect. Indeed, the presence of certain symmetries prohibits the appearance of the NH skin effect. These symmetries include PT symmetry [47] and pseudo-Hermitian symmetry [62] in the unbroken phase for any-dimensional systems as well as parity symmetry and TRS with the symmetry operator squaring to one for 1D systems [62].
@@ 5 Topological boundary states and the NH skin states
Next we turn to the coexistence of skin states and topological boundary states, and discuss how to distinguish them. Previously, we saw that the NH skin effect is protected by spectral topology, specifically the point gap topology unique to NH systems. In contrast, topological boundary states arise from wavefunction topology. If a model features both NH skin states and topological boundary states, it is a priori not obvious, how to distinguish these different types of states. We illustrate a way to make this distinction with an example.
@@
We focus on the anisotropic SSH model described by the Hamiltonian [53, 30]
$\begin{split}H=\sum_{n}\big{[}&(t_{1}+\gamma)c_{A,n}^{\dagger}c_{B,n}+(t_{1}-\gamma)c_{B,n}^{\dagger}c_{A,n}\\[-5.0pt] &+t_{2}(c_{A,n+1}^{\dagger}c_{B,n}+c_{B,n+1}^{\dagger}c_{A,n})\big{]},\end{split}$ (4)
@@
where $c_{\alpha,n}^{\dagger}$ $(c_{\alpha,n})$ creates (annihilates) a state on sublattice $\alpha$ in unit cell $n$, $t_{1}$ ($t_{2}$) is the nearest-neighbor hopping parameters inside (between) unit cells, and $\gamma$ makes the system NH by changing the magnitude of the hopping to the right with respect to the hopping to the left. This model exhibits sublattice symmetry, which allows for the explicit construction of the topological end states via a destructive interference argument [53]. Such end states are exponentially localized, and only have weight on one of the sublattices, as shown in Fig. 2(a) for a right eigenstate localized to the left boundary. Due to the anisotropy factor $\gamma$, all the other states pile up on one of the boundaries, cf. Fig. 2(b), thus displaying the NH skin effect. Let us now answer the questions as to how to distinguish the topological boundary states from the skin states.
@@
The answer to this question lies in the biorthogonal properties of these states. While the skin states behave as bulk states when studying their biorthogonal localization, as we already discussed for the Hatano-Nelson model, topological boundary states will remain localized to the boundary, as shown in Fig. 2(c). We note that situations may also arise in which the right eigenstate of a boundary state solution is localized to the opposite boundary as compared to the left eigenstate, as also mentioned in Ref. 37 in the context of topological sensors discussed below. In this case, the biorthogonal product of the right and left topological boundary state will still show a stronger signature to one of the boundaries, while preserving its general profile.
@@
As an interesting side note, we remark that this model breaks the cBBC [53, 30], where a topological invariant defined from the Bloch Hamiltonian predicts the existence of modes on the boundaries. The previously mentioned GBZ approach provides a route towards restoring the cBBC [30, 45, 44], whereas an alternative approach based on the biorthogonal properties of the boundary modes was proposed in Ref. 53. This breakdown of the cBBC first noticed in Ref. 63 is a common feature of “conventional” topological models featuring the NH skin effect [47], and can be intuitively understood from the fact due a discrepancy between the PBC and OBC spectra, gap closings indicating topological phase transitions generally occur for different parameter values.
@@ 6 Skin effect in higher dimensions
So far we have only considered the appearance of skin states in 1D systems. In higher dimension, there is not yet an all-encompassing theory. Efforts in this direction have been made very recently with the development of the amoeba formalism [36]. The amoeba, which is a mathematical region in $\mathbb{R}^{d}$ having holes and narrowing tentacles extending to infinity, is a very recent addition to algebraic geometry resembling its biological prototype [64]. In particular, the formalism relies on the principle that all algebraic-geometric information has to be encoded in the characteristic polynomials of the NH model described by the Bloch Hamiltonian. The amoeba formulation, combined with its dual Newton polygon formalism [65] and tropical geometric framework [66], provides a unified approach to characterize the NH skin effect [67, 68, 36, 69, 70].
Let us discuss this with a simple example of a single-band model in two dimensions (2D) [36], shown in Fig. 3(a). Its Hamiltonian reads
$H$ $=\sum_{n,m}\left[J_{L}\left(c^{\dagger}_{n,m}c_{n+1,m}+c^{\dagger}_{n,m+1}c_{n,m}\right)\right.$ (5)
$+J_{R}\left(c^{\dagger}_{n+1,m}c_{n,m}+c^{\dagger}_{n,m}c_{n,m+1}\right)$
$+t^{\prime}\left.\left(c^{\dagger}_{n,m}c_{n+1,m+1}+c^{\dagger}_{n,m+1}c_{n+1,m}+H.c.\right)\right],$
where $c^{\dagger}_{n,m}$ ($c_{n,m}$) creates (annihilates) an excitation on site $(n,m)$ in the 2D lattice, and $J_{L}$ and $J_{R}$ ($t^{\prime}$) are the (next-)nearest neighbor hoppings. To determine the Bloch Hamiltonian, one Fourier transforms to introduce the two Bloch momenta $k_{x}$ and $k_{y}$. Likewise to the GBZ theory, one analytically continues $k_{x,y}$, and replaces the plane waves $(e^{ik_{x}},e^{ik_{y}})$ by the complex functions $(\beta_{y},\beta_{y})$ to find the non-Bloch Hamiltonian $H(\beta_{x},\beta_{y})$. Then, the key lies in the solution of its characteristic equation
@@
Without solving this polynomial fully for $\beta_{x}$ and $\beta_{y}$, one can express $\beta_{y}(E)$ as function of $\beta_{x}(E)$ as $\beta_{y,a}(\beta_{x},E)$, where $a$ labels the different solutions in $\beta_{x}$, in our example $a=1,2$. Then, by plotting $\log|\beta_{x}(E)|$ against $\log|\beta_{y,a}(E)|$ for any $a$ and $E$ we get an amoeba, depicted in Fig. 3(b). If any amoeba for any $a$ contains no central hole, $E$ belongs to the OBC spectrum, and one can construct the full OBC spectrum in this fashion. The associated Ronkin function then provides insight into the skin state localization lengths [36]. In 1D, the construction of the amoeba reduces to the construction of the GBZ.
As a side note, we remark that the NH skin effect may also appear on higher-order boundaries in higher-dimensional models [29, 71, 72, 73], and has also been observed in higher-dimensional non-periodic systems such as quasicrystals and amorphous networks as well as fractals lattices [74].
Figure 3: Amoeba formulation in 2D. (a) Illustration of the 2D single-band model with asymmetric hopping in Eq. (5). (b) The plots depict 2D amoeba, where the points $(\log|\beta_{x}|,\log|\beta_{y}|)$ satisfy the characteristic equation in Eq. (6). A hole appears in the amoeba for energies outside the OBC spectrum, while no hole is present for energies within the OBC spectrum. The disappearance of the hole in the amoeba provides crucial information about the GBZ in the 2D system.
@@ 7 Skin effect in many-body systems
Next, we discuss the NH skin effect in the many-body context, which is an active topic of current research with significant differences between fermionic and bosonic systems. The many-body scenario is more intricate in the fermionic case due to the interplay of non-orthogonal eigenstructures and Pauli’s exclusion principle [20]. Furthermore, the fermionic repulsion significantly alters the occupied orbitals, ensuring that no more than one fermion, or one hard-core boson, can occupy each physical site. Consequently, the exponential localization of all fermions at a boundary is impossible [18, 19]. In Ref. 20, it is shown that the many-body skin effect results from an imbalance in the density distribution rather than just a sum of exponential orbitals due to the exclusion principle. The degree of asymmetry in this distribution quantifies the exponential localization of the skin modes as the system size increases. In Ref. 22, a general criterion for the appearance of the NH skin effect in many-body systems, known as the Fock-space skin effect, is discussed. Beyond the Pauli principle, interactions enrich the phenomenology of the skin effect. For instance, attractive interactions may cause clustering and thus localization, while repulsive interactions promote delocalization. In contrast to fermionic models, nonreciprocal bosonic models with interacting bosons lead to particle accumulation at the edges, forming a skin superfluid state along with distinct Mott-insulating regimes [75]. Ref. 21 introduces the NH Mott skin effect in a bosonic chain with spin degrees of freedom, emphasizing the interaction between strong correlations and NH point-gap topology.
@@
Turning to the properties of these systems, it has been shown that the skin effect suppresses entanglement propagation hindering thermalization [76]. This leads to an area law for the entanglement entropy of the non-equilibrium steady state [77], in contrast to a volume law in Hermitian systems. Many more interesting effects are predicted in the many-body context, such as skin clustering leading to Hilbert space fragmentation [78], many-body localization [79], multifractality [80], and multipole skin effect [81].
8 Experimental realizations and applications
The non-Hermitian skin effect has been observed in a plethora of classical experimental platforms, such as in coupled fiber loops [9], mechanical metamaterials [11], topoelectric circuits [12], quantum walks [10], acoustic metamaterials [13], and nano-optomechanical networks [40]. Signatures of the skin effect have also been seen in a few quantum setups such as in cold atoms [23] and in a multi-termal quantum Hall device [82]. While the skin effect can only be observed by measuring the right or left eigenstates separately, there are observables that crucially depend on the interplay between both. Schomerus shows in the context of non-reciprocal metamaterials that while the dynamical response of such systems gives direct information on the right and left eigenvectors separately, the overall sensitivity is governed by the biorthogonal properties of these eigenmodes [83].
@@ 8 Experimental realizations and applications
Beyond observing the existence of skin states, some studies focus on harnessing this effect for real world applications. For example, the aforementioned experiment in fiber loops proposes a topological funnel for light [9]. In this setup, two anisotropic 1D chains with opposite winding numbers are coupled to each other so that the skin effect in both chains conspires to act as a funnel resulting in an accumulation of all excitations at the interface. Another proposal to employ the NH skin effect is in the context of NH topological sensors [37]. Here it is shown that coupling the ends of the anisotropic SSH model, cf. Eq. (4), results in an exponential splitting of the eigenvalues, such that perturbing this coupling leads to an exponentially strong response. A crucial ingredient in this context is the presence of a single topological end state, whose right and left eigenvectors localize on opposite ends. This mechanism is exploited in the experimental realization of the NH topological ohmmeter in Ref. 38. The skin effect has also been related to directional amplification in driven-dissipative cavity arrays [39, 34]. Here, a signal entering a NH chain, similar to the Hatano-Nelson model in Eq. (1), at one end gets exponential amplified towards the other end, while it gets exponential suppressed in the other direction. This principle has been experimentally verified in a nano-optomechanical network in Ref. 40.
9 Summary and Outlook
In summary, we systematically explored the underlying mechanisms driving the NH skin effect demonstrating how it fundamentally shapes NH phases. We started by discussing several well-known features of the NH skin effect before shifting our focus to what we believe are some of the most interesting properties associated with it. After discussing how non-normality and non-reciprocity are necessary ingredients to generate a NH skin effect in NH systems, we continued by exploring the general properties of the NH skin effect, including its spectral characteristics and topological protection. Additionally, we reviewed how skin modes can be distinguished from topological boundary modes by analyzing their biorthogonal footprint. The NH skin effect in higher dimensions was discussed focusing on the recent amoeba formalism. Lastly, we explored the NH skin effect in many-body scenarios, and surveyed the experimental advances and applications.
We end this perspective with an outlook. While the NH skin effect and non-Bloch theory are well understood in 1D systems, a rigorous theoretical framework is needed for higher dimensions. Exploring connections to modern mathematics [36, 68], including K-theory and symmetry classifications, offers promising avenues for uncovering and engineering NH skin effects in higher-dimensional systems. At the same time, the interplay between interactions in NH topological phases and skin modes still needs to be fully understood. Recent simulations on skin clustering in strongly interacting bosonic systems, highlighting the fragmented Hilbert space [84], reveal new connections to phenomena like the eigenstate thermalization hypothesis and quantum scars. Extensions of the NH skin effect in 2D correlated fermion systems open new avenues for exploring many-body skin effects in higher dimensions [85]. Additionally, the impact of time-periodic driving on NH skin effects in single-particle and interacting many-body systems presents a fascinating frontier for future research [86, 87]. Another important aspect is to understand the time dynamics and steady state nature of such NH skin modes in open quantum systems [14]. Also, we want to highlight that the introduction of nonlinearities enrich the phenomenology of the skin effect [88], leading to modified localization properties, such as trap-skin states [89], and in general phenomena beyond established topological theories [90, 91].

### v2

@@ 4.1 Beyond spectral winding numbers: Symmetries
There are models, which feature a NH skin effect, while having a zero spectral winding number. In these cases, it is possible to find alternative spectral invariants. For example, the presence of so-called time-reversal dagger symmetry (TRS[60] with the symmetry operator squaring to minus one results in doubly degenerate loops in the PBC spectrum with opposite winding numbers. As a result, half of the skin states accumulate on one boundary and the other half on the other. The spectral invariant is a $\mathbb{Z}_{2}$ invariant, and this type of skin effect is referred to as the $\mathbb{Z}_{2}$-skin effect [30]. In general, symmetries play an important role in the context of the NH skin effect. Indeed, the presence of certain symmetries prohibits the appearance of the NH skin effect. These symmetries include parity-time symmetry [43] and pseudo-Hermitian symmetry [60] in the unbroken phase for any-dimensional systems as well as parity symmetry and TRS with the symmetry operator squaring to one for 1D systems [60].
@@ 5 Topological boundary states and the NH skin states
Next we turn to the coexistence of skin states and topological boundary states, and discuss how to distinguish them, which we believe is important to understand. Previously, we saw that the NH skin effect is protected by spectral topology, specifically the point gap topology unique to NH systems. In contrast, topological boundary states arise from wavefunction topology. If a model features both NH skin states and topological boundary states, it is a priori not obvious how to distinguish these different types of states.
@@
To illustrate this distinction, we focus on the anisotropic SSH model described by the Hamiltonian [49, 32]
$H=\sum_{n}\big{[}(t_{1}+\gamma)c_{A,n}^{\dagger}c_{B,n}+(t_{1}-\gamma)c_{B,n}^{\dagger}c_{A,n}\\[-5.0pt] +t_{2}(c_{A,n+1}^{\dagger}c_{B,n}+c_{B,n+1}^{\dagger}c_{A,n})\big{]},$ (4)
@@
where $c_{\alpha,n}^{\dagger}$ $(c_{\alpha,n})$ creates (annihilates) an excitation on sublattice $\alpha=A,B$ in unit cell $n$, $t_{1}$ ($t_{2}$) is the NN hopping parameters inside (between) unit cells, and $\gamma$ makes the system NH by changing the magnitude of the hopping to the right with respect to the hopping to the left, i.e., it makes the hoppings asymmetric. This model exhibits sublattice symmetry, which allows for the explicit construction of the topological end states via a destructive interference argument [49]. Such end states are also exponentially localized, however, they only have weight on one of the sublattices, as shown in Fig. 2(a) for a right eigenstate localized to the left boundary. Due to the anisotropy factor $\gamma$, all the other states pile up on one of the boundaries, cf. Fig. 2(b), thus displaying the NH skin effect. Let us now answer the question as to how to distinguish the topological boundary states from the skin states.
@@
The answer to this question lies in the biorthogonal properties of these states. While the skin states behave as bulk states when studying their biorthogonal localization, as we already discussed before, topological boundary states will remain localized to the boundary, as shown in Fig. 2(c). We note that situations may also arise in which the right eigenstate of a boundary-state solution is localized to the opposite boundary as compared to the left eigenstate, as also mentioned in Ref. 61 in the context of topological sensors discussed below. In this case, the biorthogonal product of the right and left topological boundary state will still show a stronger signature to one of the boundaries, while preserving its general profile.
@@
We remark that this model breaks the cBBC [49, 32], where a topological invariant defined from the Bloch Hamiltonian predicts the existence of modes on the boundaries. The previously mentioned GBZ approach provides a route towards restoring the cBBC [32, 41, 40], whereas an alternative approach based on the biorthogonal properties of the boundary modes was proposed in Ref. 49. This breakdown of the cBBC, first noticed in Ref. 56, is a common feature of topological models featuring the NH skin effect [43], and can be intuitively understood from the fact that a discrepancy between the PBC and OBC spectra means gap closings indicating topological phase transitions generally occur for different parameter values.
@@ 6 Skin effect in higher dimensions
So far we have only considered the appearance of skin states in 1D systems. In higher dimension, there is not yet an all-encompassing theory, which we believe to be an important future direction of study. Efforts in this area have been made very recently with the development of the amoeba formalism [62], the dual Newton polygon formalism [63], and the tropical geometric framework [64] to characterize the NH skin effect [65, 66, 62, 67, 68].
To highlight these new developments, let us discuss the amoeba formalism with a simple example of a single-band model in two dimensions (2D) [62], shown in Fig. 3(a). Its Hamiltonian reads
$H$ $=\sum_{n,m}\left[J_{L}\left(c^{\dagger}_{n,m}c_{n+1,m}+c^{\dagger}_{n,m+1}c_{n,m}\right)\right.$
$\quad\ +J_{R}\left(c^{\dagger}_{n+1,m}c_{n,m}+c^{\dagger}_{n,m}c_{n,m+1}\right)$
[…]
where $c^{\dagger}_{n,m}$ ($c_{n,m}$) creates (annihilates) an excitation on site $(n,m)$ in the 2D lattice, and $J_{L}$ and $J_{R}$ ($t$) are the (next-)nearest neighbor hoppings. To determine the Bloch Hamiltonian, one Fourier transforms to introduce the two Bloch momenta $k_{x}$ and $k_{y}$. Likewise to the GBZ theory, one analytically continues the Bloch momenta, i.e., one replaces the plane waves $(e^{ik_{x}},e^{ik_{y}})$ by the complex functions $(\beta_{x},\beta_{y})$ to find the non-Bloch Hamiltonian $H(\beta_{x},\beta_{y})$. Then, the formalism relies on the principle that all algebraic-geometric information has to be encoded in its characteristic equation
@@
An amoeba is a region in $\mathbb{R}^{d}$ solving this equation ($d=2$ in our example), resembling its biological prototype [69], having holes and narrowing tentacles extending to infinity.
In the example, we can construct the amoebas without solving this polynomial fully for $\beta_{x}$ and $\beta_{y}$. Expressing $\beta_{y}(E)$ as a function of $\beta_{x}(E)$ as $\beta_{y,A}(\beta_{x},E)$, where $A=1,2$ labels the different solutions in $\beta_{x}$, we can plot $\log|\beta_{x}(E)|$ against $\log|\beta_{y,A}(E)|$ for any $A$ and $E$ to get amoebas, depicted in Fig. 3(b,c). If any amoeba for any $A$ contains no central hole, $E$ belongs to the OBC spectrum, and one can construct the full OBC spectrum in this fashion. The associated Ronkin function then provides insight into the skin state localization lengths [62]. In 1D, the construction of the amoeba reduces to the construction of the GBZ.
We also remark that the NH skin effect may also appear on higher-order boundaries in higher-dimensional models [31, 70, 71, 72], and has also been observed in higher-dimensional non-periodic systems such as quasicrystals and amorphous networks as well as fractals lattices [73].
Figure 3: Amoeba formulation in 2D. (a) Illustration of the 2D single-band model with asymmetric hoppings, Eq. (6). (b,c) The plots depict 2D amoeba, where the points $(\log|\beta_{x}|,\log|\beta_{y}|)$ satisfy the characteristic equation Eq. (6). A hole appears in the amoeba for energies outside the OBC spectrum in (b), while in (c), no hole is present for energies within the OBC spectrum. We set $J_{R}=4/5,J_{R}=6/5$, and $t=1/2$.
@@ 7 Skin effect in many-body systems
Another active topic of current research we find important and want to highlight is the NH skin effect in the context of many-body physics. Here, fermionic and bosonic systems show significant differences. For example, the fermionic case is quite intricate due to the interplay of non-orthogonal eigenstructures and Pauli’s exclusion principle [21]. Also, fermionic repulsion significantly alters the occupied orbitals, ensuring that no more than one fermion, or one hard-core boson, can occupy each physical site. Consequently, the exponential localization of all fermions at a boundary is impossible [19, 20]. In Ref. 21, it is shown that the many-body skin effect results from an imbalance in the density distribution rather than just a sum of exponential orbitals due to the exclusion principle. The degree of asymmetry in this distribution quantifies the exponential localization of the skin modes as the system size increases. In Ref. 23, a general criterion for the appearance of the NH skin effect in many-body systems, known as the Fock-space skin effect, is discussed. Beyond the Pauli principle, interactions enrich the phenomenology of the skin effect. For instance, attractive interactions may cause clustering and thus localization, while repulsive interactions promote delocalization. In contrast to fermionic models, nonreciprocal bosonic models with interacting bosons lead to particle accumulation at the edges, forming a skin superfluid state along with distinct Mott-insulating regimes [74]. Ref. 22 introduces the NH Mott skin effect in a bosonic chain with spin degrees of freedom, emphasizing the interaction between strong correlations and NH point-gap topology.
@@
Turning to the properties of these systems, it has been shown that the skin effect suppresses entanglement propagation hindering thermalization [75]. This leads to an area law for the entanglement entropy of the non-equilibrium steady state [76], in contrast to a volume law in Hermitian systems. More interesting effects are predicted in the many-body context, e.g., skin clustering leading to Hilbert space fragmentation [77], many-body localization [78], multifractality [79], and multipole skin effect [80].
8 Experimental signatures
The NH skin effect has been observed in a plethora of classical experimental platforms, such as in coupled fiber loops [9], mechanical metamaterials [11], topoelectric circuits [12], quantum walks [10], acoustic metamaterials [13], nano-optomechanical networks [81], and thermal diffusion [82]. A particular challenge to observe the NH skin effect is that one has to implement either asymmetric hoppings or complex phases with on-site losses. In this context, we want to highlight the setup in Ref. 83, a 2D photonic metamaterial on a chip, which allows to dynamically reconfigure imaginary gauge fields, corresponding to complex phases on the hoppings. Thus, one is able to manipulate the skin effect on the fly in a compact, versatile and scalable platform, promising future applications.
Furthermore, signatures of the skin effect have also been seen in quantum setups such as in cold atoms [24] and in a multi-terminal quantum Hall device [84]. One of such signatures is called self-acceleration, an early-time effect in the dynamics of the system [85], was observed in a quantum walk setup [86].
While the skin effect can only be directly observed by measuring the right or left eigenstates separately, it is possible to reconstruct them by measuring the system’s response. This was done in Ref. 87 in the model discussed in the context of the amoeba formulation, cf. Eq. (6). Measuring the system’s response, the Green’s function was determined and from it the full complex spectrum as well as the left and right eigenstates. In general, it has been shown that there are observables that crucially depend on the interplay between both flavors of eigenstates, such as the state skewness in the example above, or the system’s sensitivity against perturbations as pointed out in this important paper by Schomerus [88].
9 Applications
@@
Beyond observing the existence and effects of skin states, an important development we want to highlight is the harnessing of this effect for real world applications. For example, an experiment using fiber loops proposes a topological funnel for light [9]. In this setup, two anisotropic 1D chains with opposite winding numbers are coupled to each other so that the skin effect in both chains conspires to act as a funnel resulting in an accumulation of all excitations at the interface. Another proposal to employ the NH skin effect is in the context of NH topological sensors [61]. Here it is shown that coupling the ends of the anisotropic SSH model, cf. Eq. (4), results in an exponential splitting of the eigenvalues, such that perturbing this coupling leads to an exponentially strong response. A crucial ingredient in this context is the presence of a single topological end state, whose right and left eigenvectors localize on opposite ends. This mechanism is exploited in the experimental realization of the NH topological ohmmeter in Ref. 89. Other NH sensing protocols exist, such as in Ref. 90.
The skin effect has also been related to directional amplification [91, 92, 53]. Here, a signal entering a NH chain at one end gets exponential amplified towards the other end, while it gets exponential suppressed in the other direction. This principle has been experimentally verified in the bosonic Kitaev chain [93], a model similar to the Hatano-Nelson model in Eq. (1), in a nano-optomechanical network [81] and in a quantum simulator in Ref. [94].
10 Outlook
Let us end with an outlook. While the NH skin effect is well understood in 1D systems, a complete theoretical framework is needed for higher dimensions, where we discussed the amoeba formalism as one important cornerstone to tackle such problems. Exploring connections to modern mathematics, including K-theory and symmetry classifications, offers promising avenues for uncovering and engineering the NH skin effect in higher-dimensional systems.
At the same time, the interplay between interactions in NH topological phases and skin modes still needs to be fully understood. For example, recent simulations on skin clustering in strongly interacting bosonic systems highlight the fragmented Hilbert space [95] and reveal new connections to phenomena like the eigenstate thermalization hypothesis and quantum scars. Also, going to higher dimensions in correlated fermionic systems opens new avenues to explore many-body skin effects [96].
Beyond what was touched on in the main text, we see great prospects in analyzing the impact of time-periodic driving, i.e., Floquet engineering, in single-particle and interacting many-body systems [97, 98]. Also time dynamics and the steady state nature of NH skin modes in open quantum systems promise new insights [14]. Finally, we see the introduction of nonlinearities [99] as a great source for novel phenomenology related to the skin effect, leading to modified localization properties, such as trap-skin states [100], and in general, phenomena beyond established topological theories [101, 102].